# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

### v2

@@ Appendix D Changes in the Parts Without Regularization
Figure 6: Ratio of the sum of absolute values of unregularized parameters after LLaMA2-7B regularization to the sum of absolute values of the corresponding parameters before regularization. For example, regularization is not applied to the first 75% of rows in $W_{Q}$, and the absolute sum of the rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Increased Ratios. This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the unregularized columns is compared to the original sum.
Appendix E Performance of DReSS under Different Pruning Ratios and Datasets
E.1 The Perplexity of DReSS under Different Pruning Ratios and Datasets
@@
Model Pruning Ratio WikiText-2 Alpaca PTB C4 OPT-2.7B Dense 12.46 11.64 17.97 14.32 10% 12.48 11.71 18.16 14.71 20% 12.51 11.93 19.45 15.42 30% 14.49 12.88 23.58 18.93 40% 18.92 15.46 31.30 24.33 50% 24.57 21.37 45.32 32.15 60% 33.83 31.22 58.73 45.56 OPT-6.7B Dense 10.85 10.27 15.77 12.71 10% 10.91 10.45 16.05 13.18 20% 11.02 10.83 17.58 14.45 30% 12.32 11.46 19.65 16.68 40% 14.26 12.71 25.52 20.94 50% 19.63 15.66 33.78 28.22 60% 28.75 21.39 47.29 39.47 OPT-13B Dense 10.12 9.46 14.52 12.06 10% 10.26 9.65 14.88 12.37 20% 10.47 9.97 15.64 13.15 30% 10.99 10.56 18.81 15.63 40% 11.62 11.25 23.07 19.55 50% 16.15 14.68 29.26 26.21 60% 27.63 20.12 38.59 35.58 LLaMA2-7B Dense 5.47 5.25 7.92 7.26 10% 5.54 5.29 8.06 7.34 20% 5.63 5.37 8.29 7.79 30% 7.39 7.32 9.13 8.46 40% 9.62 8.62 12.37 10.56 50% 13.77 12.59 18.85 15.18 60% 24.38 20.25 29.34 27.37 LLaMA2-13B Dense 4.88 4.63 7.16 6.73 10% 4.94 4.69 7.23 6.96 20% 5.08 4.83 7.61 7.53 30% 5.61 5.42 8.33 8.14 40% 6.25 6.14 9.27 9.05 50% 7.59 7.28 11.58 10.88 60% 12.77 11.78 15.16 13.62
E.2 The Accuracy of DReSS under Different Pruning Ratios on Zero-shot Tasks
@@
Model Pruning Ratio PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39 10% 70.38 60.12 51.79 52.46 29.78 52.91 20% 69.96 59.47 50.45 51.87 28.62 52.07 30% 64.52 58.76 49.25 50.23 27.52 50.06 40% 61.32 56.27 47.39 48.26 25.57 47.76 50% 59.38 53.46 44.63 46.19 21.66 45.06 60% 52.99 48.74 40.03 43.35 16.25 40.27 OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70 10% 75.89 64.61 64.69 58.53 33.47 59.44 20% 75.12 64.23 62.56 57.34 32.95 58.44 30% 72.52 62.63 58.27 54.48 29.99 55.58 40% 67.37 58.59 52.12 49.38 26.87 50.87 50% 61.48 55.62 46.46 47.68 22.97 46.84 60% 55.73 51.52 43.38 45.85 18.62 43.02 OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79 10% 75.46 64.69 68.56 61.79 35.58 61.22 20% 74.78 64.61 67.25 61.66 35.43 60.75 30% 72.62 63.26 65.69 59.64 32.57 58.76 40% 68.67 61.49 62.74 55.98 29.26 55.63 50% 62.19 56.46 58.55 52.15 24.53 50.78 60% 57.43 52.72 51.23 47.62 21.05 46.01 LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00 10% 77.38 68.16 71.28 69.26 43.73 65.96 20% 76.42 67.35 68.26 66.73 41.68 64.09 30% 72.29 64.87 58.53 63.48 39.27 59.69 40% 68.66 61.49 55.42 58.61 36.52 56.14 50% 61.32 55.68 51.79 52.35 31.55 50.54 60% 56.45 51.79 48.96 48.98 27.26 46.69 LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76 10% 79.35 72.15 76.42 76.92 48.57 70.68 20% 78.27 71.98 73.20 75.43 47.62 69.30 30% 75.49 69.73 71.54 72.87 45.41 67.00 40% 73.56 64.46 67.75 68.45 41.28 63.10 50% 68.73 59.82 62.48 60.12 36.14 57.46 60% 62.25 56.27 59.93 53.37 29.77 52.32
Appendix F More Details of the Dependency on Calibration Dataset
When using Alpaca, WikiText-2, C4, and PTB as calibration and RFT data, the perplexity of various methods on Alpaca, C4, and PTB is shown as follows:

### v3

@@ Appendix F Changes in the Parts Without Regularization
Figure 6: Ratio of the sum of absolute values of unregularized parameters after LLaMA2-7B regularization to the sum of absolute values of the corresponding parameters before regularization. For example, regularization is not applied to the first 75% of rows in $\mathbf{W_{Q}}$, and the absolute sum of the rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Increased Ratios. This is applied similarly to other matrices such as $\mathbf{W_{K},W_{V},W_{o},W_{gate},W_{up},W_{down}}$, and for layer activations, the sum of the unregularized columns is compared to the original sum.
Appendix G Performance of DReSS under Different Pruning Ratios and Datasets
G.1 The Perplexity of DReSS under Different Pruning Ratios and Datasets
@@
Model Pruning Ratio WikiText-2 Alpaca PTB C4
OPT-2.7B Dense 12.46 11.64 17.97 14.32
10% 12.48 11.71 18.16 14.71
20% 12.50 11.93 19.45 15.42
30% 14.49 12.88 23.58 18.93
40% 18.92 15.46 31.30 24.33
50% 24.57 21.37 45.32 32.15
60% 33.83 31.22 58.73 45.56
OPT-6.7B Dense 10.85 10.27 15.77 12.71
10% 10.91 10.45 16.05 13.18
20% 11.02 10.83 17.58 14.45
30% 12.32 11.46 19.65 16.68
40% 14.26 12.71 25.52 20.94
50% 19.63 15.66 33.78 28.22
60% 28.75 21.39 47.29 39.47
OPT-13B Dense 10.12 9.46 14.52 12.06
10% 10.22 9.65 14.88 12.37
20% 10.31 9.97 15.64 13.15
30% 10.99 10.56 18.81 15.63
40% 11.62 11.25 23.07 19.55
50% 13.85 13.23 29.26 26.21
60% 27.63 20.12 38.59 35.58
LLaMA2-7B Dense 5.47 5.25 7.92 7.26
10% 5.54 5.29 8.06 7.34
20% 5.63 5.37 8.29 7.79
30% 7.39 7.32 9.13 8.46
40% 9.62 8.62 12.37 10.56
50% 13.77 12.59 18.85 15.18
60% 24.38 20.25 29.34 27.37
LLaMA2-13B Dense 4.88 4.63 7.16 6.73
10% 4.94 4.69 7.23 6.96
20% 5.08 4.83 7.61 7.53
30% 5.61 5.42 8.33 8.14
40% 6.25 6.14 9.27 9.05
50% 7.59 7.28 11.58 10.88
60% 12.77 11.78 15.16 13.62
@@
G.2 The Accuracy of DReSS under Different Pruning Ratios on Zero-shot Tasks
@@
Model Pruning Ratio PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc
OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39
10% 70.38 60.12 51.79 52.46 29.78 52.91
20% 69.96 59.47 50.45 51.87 28.62 52.07
30% 64.52 58.76 48.25 50.23 27.52 49.86
40% 61.32 56.27 47.39 48.26 25.57 47.76
50% 59.38 53.46 44.63 46.19 21.66 45.06
60% 52.99 48.74 40.03 43.35 16.25 40.27
OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70
10% 75.89 64.61 64.69 58.53 33.47 59.44
20% 75.12 64.23 62.56 57.34 32.95 58.44
30% 72.52 62.63 58.27 54.48 29.99 55.58
40% 67.37 58.59 52.12 49.38 26.87 50.87
50% 61.48 55.62 46.46 47.68 22.97 46.84
60% 55.73 51.52 43.38 45.85 18.62 43.02
OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79
10% 75.46 64.69 68.56 61.79 35.58 61.22
20% 74.78 64.61 67.25 61.66 35.43 60.75
30% 72.62 63.26 65.69 59.64 32.57 58.76
40% 68.67 61.49 62.74 55.98 29.26 55.63
50% 62.19 56.46 58.55 52.15 24.53 50.78
60% 57.43 52.72 51.23 47.62 21.05 46.01
LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00
10% 77.38 68.16 71.28 69.26 43.73 65.96
20% 76.42 67.35 68.26 66.73 41.68 64.09
30% 72.29 64.87 58.53 63.48 39.27 59.69
40% 68.66 61.49 55.42 58.61 36.52 56.14
50% 61.32 55.68 51.79 52.35 31.55 50.54
60% 56.45 51.79 48.96 48.98 27.26 46.69
LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76
10% 79.35 72.15 77.42 76.92 48.57 70.88
20% 78.27 71.98 76.89 75.43 47.62 70.04
30% 75.49 69.73 73.54 72.87 45.41 67.41
40% 73.56 64.46 67.75 68.45 41.28 63.10
50% 68.73 59.82 62.48 60.12 36.14 57.46
60% 62.25 56.27 59.93 53.37 29.77 52.32
@@
Appendix H Dependency on Calibration Dataset
Since DReSS relies on data-driven regularization, we investigate its dataset dependency. We evaluated perplexity of four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4. For fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2. As shown in Figure 7, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness. When using Alpaca, WikiText-2, C4, and PTB as calibration and RFT data, the perplexity of various methods on WikiText-2, Alpaca, C4, and PTB is shown as follows:
Figure 7: Comparison of perplexity on Wikitext-2 using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
@@ Appendix H Dependency on Calibration Dataset
Appendix I Future Work and Limitations
The “regularize-then-prune-then-finetune” paradigm proposed by DReSS is generalizable. In this paper, we only explore pruning at the channel level (i.e., rows or columns of parameter matrices). In future work, we plan to extend the pruning units to entire transformer layers. Using transformer layer as pruning unit is expected to yield more significant acceleration.